OpenStudy (anonymous):

help Which equation does the graph of the systems of equations solve? 2 linear graphs. They intersect at 0,3

OpenStudy (anonymous):

Answer choices A) −2x + 3 = 8x − 3 B) −2x + 4 = −8x + 4 C) 2x + 4 = 8x − 4 D) 2x + 3 = −8x + 3

zepdrix (zepdrix):

Notice that they intersect at an x coordinate of 0.

zepdrix (zepdrix):

Let's look at option A. -2x + 3 = 8x - 3 `I'll add 2x to each side` 3 = 10x - 3 `I'll add 3 to each side` 6 = 10x Hmm this is NOT going to lead to a solution of x=0, which is what we want.

zepdrix (zepdrix):

Do you think you can try the other options? :o Solve for x in each case.

OpenStudy (anonymous):

yes let me try

zepdrix (zepdrix):

Oh that was a sloppy method of doing that actually :\ That's going to only narrow it down to two options. I guess a better way to do it would be... plug in x=0, and see which equation gives you 3 = 3.

OpenStudy (anonymous):

okay let me try doing it that way can you show me an example please

zepdrix (zepdrix):

For Option A, \(\large\rm -2x + 3 = 8x - 3\) We know that the point of intersection is at (0,3), so I'll plug in this x coordinate of 0, and it should give us 3 as a result on both sides, \(\large\rm -2(0)+3=8(0)-3\) Which simplifies to \(\large\rm 0+3=0-3\) And since \(\large\rm 3\ne-3\), this is not the correct option for us.

OpenStudy (anonymous):

okay I'll try it for the other options, give me few seconds please thank you

OpenStudy (anonymous):

it would be option D

zepdrix (zepdrix):

Yay good job \c:/ Notice that using the first method I explained.. Would have given you option B as a correct choice, so that was no beuno. This worked out much better! :)

OpenStudy (anonymous):

yes that's correct, thank you very much